A mathematics test was given to 50girls and 75 boys.The girl made an average grade of 76 with an SD of 6 and the boys made an average grade of 82 with an SD of 2. Test whether there is any difference between the performance of boys and girls
1
Expert's answer
2021-07-19T05:51:39-0400

"n_1 = 50 \\

n_2=75 \\

\bar{x_1}=75 \\

s_1=6 \\

\bar{x_2}= 82 \\

s_2=2 \\

H_0: \mu_1 = \mu_2 \\

H_1: \mu_1 ≠ \mu_2"

Test-statistic:

"Z = \frac{\bar{x_1} -\bar{x_2}}{\sqrt{ (s_1^2/n_1)+(s^2_2/n_2) }} \\

Z= \frac{75-82}{\sqrt{ (6^2/50) +(2^2/75) }} = -7.96"

Let use 5 % significant level.

Two-tailed test.

Reject H0 if Z≤ -1.96 or Z≥1.96.

"Z= -7.96 < Z_{crit}= -1.96"

Reject H0.

There is enough evidence to conclude that there is any difference between the performance of boys and girls at 5% significant level.